from the given information, find the equation of the line both in slope-intercept form and standard form then it says the slope contains a negative one half and contains poin negative two comma five how do I do this stuff!!! can you give me an anser and an explanation?
1
Expert's answer
2014-10-16T14:36:22-0400
The given information is not accurate.
The first case. If you meant that we know points A(-1, 1/2) and B(-2,5) of a straight line, then the slope of the line is calculated by k=(yB-yA)/(xB-xA)=(5-1/2)/(-2-(-1))=4.5/(-1)=-4.5. The equation of straight line in slope-intercept form is y=kx+b, hence knowing that point B is a point of this line and k=-4.5, conclude that b=yB-kxB, b=4-(-4.5)(-2)=4-9=-5. Finally, equation of straight line in slope-intercept form is y=-4.5x-5. Next, convert it to standard form 4.5x+y=-5.
The second case. If you meant that we know slope k=-1/2 and point B(-2,5), then equation of straight line is
y-yB=k(x-xB), y-5=-1/2*(x-(-2)), y-5=-1/2*(x+2), y-5=-1/2*x-1, hence y=-x/2+4 is the equation of straight line in slope-intercept form. Next, convert it to standard form x/2+y=4.